OpenStudy (anonymous):

Solve these equations a) 3(x+2)=12 b) 3(x+2)= 2x c) 3(x+2)=x-18 d) 3(x+2) = -15

OpenStudy (anonymous):

c) 3(x+2)=2x*

OpenStudy (anonymous):

b) 3(x+2)=x-18

OpenStudy (anonymous):

The plan is to do what is needed to get x by itself as x = etc. for 3(x+2) = 12 divide by 3 x+2 = 4 subtract 2 x = 2

OpenStudy (yanasidlinskiy):